A bag contains 8 green balls and 4 black balls. 3.1.1.Find the probability of drawing (in succession) 2 green balls whenthe first ball is replaced in the bag.[3]3.1.2.Find the probability of drawing (in succession) 2 green balls whenthe first ball is not replaced in the bag.[3]3.2.Three trucksserve a small town. TruckA has 50% of all the parcels delivered, TruckB has30% and TruckC has the remaining 20%. Their on-time deliveryrates are 80%, 65% and 40% respectively. If a parcel is chosen at random.3.2.1.What is the probability that it was deliveredon time?[5]3.2.2.If a parcel is deliveredon time, what is the probability that it was delivered by Truck
Answer by Boreal(15235)

with replacement it is (8/12)(8/12)=4/9 probability
without replacement, it is (8/12)(7/11)=14/33 probability.
For 100 parcels
50 are A with 80% on time, so 40 are on time
30 are B with 65% on time, so 19.5 are on time
20 are C with 40% on time, so 8 are on time
they add to 67.5 out of 100 or 67.5% probability.
The last part doesn't say which truck, so I will assume A--40/67.5 were from Truck A, so probability it was from A is 59.3%